OpenStudy (xbeenchantedx):

find the value of "v" for: v+2/v + 4/3v = 11

OpenStudy (xbeenchantedx):

howd u get that answer? i got 2/6 and it didnt check

OpenStudy (anonymous):

(v+2)/v + 4/3v = 11 (3v+6)/3v + 4/3v =11 (3v+10)/3v= 11 3v+10 = 33v 10= 30v v= 1/3

OpenStudy (anonymous):

My bad, I messed up along the way

OpenStudy (xbeenchantedx):

its still confusing. i got the same thing except the answer

OpenStudy (anonymous):

1/3 is equal to 2/6